OpenStudy (richyw):

can someone help explain this step to me? I have \[\left(\frac{E_y}{b}-\frac{E_x}{a}\cos\delta\right)^2=\left(1-\frac{E_x^2}{a^2}\right)\sin^2\delta\]and apparantly that is equivalent to\[\frac{E_y^2}{b^2}+\frac{E_x^2}{a^2}-2\frac{E_xE_y}{ab}\cos{\delta}=\sin^2{\delta}\]

OpenStudy (richyw):

when I expand out the left side though I get \[\frac{x^2 \cos ^2(d)}{a^2}-\frac{2 x y \cos (d)}{a b}+\frac{y^2}{b^2}\](I just made \(E_x=x,\; E_y=y\) anad \(\delta=d\) to save time

OpenStudy (richyw):

so my problem is how the \(\cos^2{d}\) gets out of that first term.

OpenStudy (richyw):

aha . nevermind I got it.